Towards the end of today's store Dana and I went though two seperate changes in the ZDCs and trigger as a step toward a final setup of these system durings this years run. a) mid- last week the ZDC input was split (50-50) and the gain of tubes lowered by a factor of two to minimize the effect of saturation in tubes (for mid-central collisions) and saturation in linear fan i/o. We also observe a significant +- 5mV noise on all linear fan i/o that effects the ZDC energy signals and disc. This noise has appeared since last year likely by addition of detectors and introduction of ground loops. This does mean that the pedestals of ZDC energies are quite bad. This should have been improved by adding 'filters' in the adc lines but some of the pedestal are still bad. A set of attenuators in the low-gain side has been removed. Thus expect that calibrations of both ADC and TDC should be modified follwoing todays runs (aug 31, 2001) b) The triggers were changed to have the following definitions -- trigger 4 Y.B.Rhic CLock -- trigger 1 BeamBeam AND Rhic Clock. The hope is to removed sets of events that seems to be picked up by 'reflection' pulses in the ZDCs and possible other false events. The timing of Lvl0 is very close to the previous setup by may have shifted by up to 1 nsec. We urge people looking at the data to report/discuss changes in data quality, timing boith if iomproved and if there are problems. I should be pointed out that we have yet to record a full run with this new configuration. Flemming ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov
This archive was generated by hypermail 2b30 : Fri Aug 31 2001 - 20:16:11 EDT